OpenStudy (anonymous):

6.An electric iron has a rating of 750W,220V.Calculate:(a)current passing through it and (b)Its resistance when in use. 7.Two lamps,one rated 100W,220V and other 40W ,220V are connected in parallel to 220V mains supply.Calculate the current drawn from the supply line. 8.An electric lamp is marked 100W,220V.Its used for 5hrs daily (a)Its resistance when in use. (b)Energy consumed in kWh per day.

OpenStudy (shane_b):

This should help you do the first one...when you're ready I can help you with the others:

OpenStudy (anonymous):

Thank you so much for your generous help, Shane_B! Your pie chart really helped a lot! Please help me with the others too!!

OpenStudy (shane_b):

For #7, first find the resistances of each lamp using the same method you used in #6. Since it's only two resistances you can use this formula to calculate the total resistance:\[\frac{R_1R_2}{R_1+R_2}\]Then use I=V/R to find the current. I have to run some errands but when I get back if no one has answered the last one I'll help you with it :)

OpenStudy (anonymous):

Thank you , no problem!

OpenStudy (shane_b):

For the last one part a can be done the same as the others. Part b is pretty very simple. The iron is rated at 100W which is 0.1kW and it's used for 5 hours per day. Therefore it's:\[kWh_{per~day}=0.1kW(5h)=0.5~kWh\]